Consider the following two functions:

1. f : R → R defined by f(x) = 4x − 15.

2. g : R → R defined by f(x) = 15x3.

Prove that both f and g are one-to-one correspondences.

Let f : A → B be a one-to-one correspondence. Then to each bB there corresponds a unique aA such that f(a) = b. We define f-1: B → A by

f-1(b) = the unique a such that f(a) = b.


1
Expert's answer
2021-03-24T15:19:43-0400

"\text{Let a,b}\in R"

"f(a)=4a-15"

"f(b) = 4b-15"

"\text{Assuming f(a) = f(b)}"

"\implies 4a-15=4b-15"

"\implies a=b"

"\text{Hence f is one to one correspondence since f(a) = f(b) implies a = b }"

"\text{Also let } g(a)=15a^3 \text{ and }g(b)=15b^3"

"\text{Assuming g(a) = g(b)}"

"\implies 15a^3=15b^3"

"\implies a=b"

"\text{Hence g is one to one correspondence since g(a) = g(b) implies a = b }"